53. This reaction was monitored as a function of time: A-→ B + C A plot of In[A] versus time yields a straight line with slope-0.0045/s. a. What is the value of the rate constant ()for this reaction at this a. What is the value of the rate constant (k) for this reaction at this temperature? b. Write the rate law for the reaction. c. What is the half-life? d. If the initial concentration of A is 0.250 M, what is the concentra- tion after 225 s?
Answer #1


a. if ln[A] vs Time , slope is -ve indicates first order reaction

   K = - slope

     = - (-0.0045)

     = 0.0045 s-1

b. rate = K[A]

c. half-life(T1/2) = 0.693/k = 0.693/0.0045 = 154 sec

d. K = (1/t)ln(a0/a0-x)

     0.0045 = (1/225)ln(0.25/(a0-x))

a0-x = concentration after 225 sec = 0.091 M